# Superconductivity in three-dimensional interacting doped topological insulators

### v1

@@ Abstract
Three-dimensional doped Dirac insulators foster simply connected (in both topological and trivial regimes) and annular (deep inside the topological regime) Fermi surfaces (FSs) in the normal state, and allow on-site repulsions among fermions with opposite spin ($U_{1}$) and parity ($U_{2}$) eigenvalues. From an unbiased leading-order (one-loop) renormalization group analysis, controlled by a suitable $\epsilon$ expansion, we show that this system develops strong propensity toward the nucleation of scalar $s$-wave and odd-parity pseudoscalar $p$-wave pairings, favored by repulsive $U_{1}$ and $U_{2}$ interactions, respectively, irrespective of the underlying FS topology. Our results can be pertinent for the observed superconductivity in various doped narrow gap semiconductors, and the theoretical foundation can readily be applied to investigate similar phenomenon in various doped topological materials.
@@
Introduction. From the baryonic universe to topological quantum crystals, Dirac theory offers a unified description. As such electrical and thermal topological insulators (TIs) belonging to any Altland-Zirnbauer symmetry class in any dimension ($d$) can be modeled by suitable massive Dirac Hamiltonian [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The ${\bf k}\cdot{\bf p}$ Hamiltonian for their candidate materials also assumes a similar form in the low energy and long-wavelength limit [16]. Fascinatingly, three-dimensional (3D) doped or intercalated TIs, (Cu/Nd/Sr)xBi2Se3 and Sn1-xInxTe, with a Fermi surface (FS), become a fully gapped superconductor at the low temperatures [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The paired state is possibly topological ($p$-wave) in nature [27, 28, 29, 30, 31], showing a zero bias conductance peak (ZBCP), stemming from surface Majorana fermions. It is also conceivable to induce a transition between topological $p$-wave and trivial $s$-wave (devoid of ZBCP) pairings in Sn1-xInxTe. Despite such exciting experimental findings in the last decade, the phase diagram of interacting doped TIs, fostering superconducting ground states still remains hazy.
@@
Doped Dirac insulators accommodate simply connected and annular FSs, see Fig. 1(a). In this work, we theoretically study the effects of electron-electron (e-e) interactions on the phase diagram of such systems using an unbiased leading-order renormalization group (RG) analysis, controlled by an appropriate $\epsilon$ expansion. In strongly correlated TIs, such as the ones in Kondo systems [32, 33, 34], e-e interactions can arise from Hubbardlike repulsions. In weakly correlated materials, effective e-e interactions can be mediated by optical phonons below the scale of optical frequency [35, 36, 37]. Without delving into their exact microscopic origin, here we consider all symmetry allowed e-e interactions. Then within the framework of a Hubbardlike model, composed of the on-site repulsions among fermions with opposite spin ($U_{1}$) and parity ($U_{2}$) eigenvalues, we show that 3D correlated doped TIs feature an intriguing confluence of odd-parity pseudoscalar topological pairing and trivial scalar $s$-wave pairing, respectively favored by $U_{2}$ and $U_{1}$, irrespective of the underlying FS geometry. See Fig. 1(b)-(d).
@@
Figure 1: (a) Fermi surface (FS) geometries for various chemical doping (dashed colored lines) in a Dirac insulator in the topological ($m<0$) and trivial ($m>0$) regimes. Here we assume $b>0$. A quantum phase transition between them takes place at $m=0$, around which the FS is simply connected (red spheres). Deep inside the topological regime ($m<-1/(2b)$), the system supports annular FS (concentric blue spheres). Black curves show energy eigenvalues (vertical direction) along a momentum (horizontal direction). Phase diagram in the $(U_{1},U_{2})$ plane for (b) $\mu=0.685$, $m=0.7$, $b=1.0$ (yielding annular FS in doped TI), (c) $\mu=0.730$, $m=-0.7$, $b=1.0$ (yielding simply connected FS in doped TI), (d) $\mu=0.180$, $m=0.1$, $b=0.1$ (yielding simply connected FS in doped trivial insulator) and $t=10^{-6}$ (mimics zero temperature). Color map in panels (b)-(d) indicates the transition temperature ($t_{c}$). White regions mark a Fermi liquid without any ordering. The phase diagrams display odd-parity topological $p$-wave and trivial $s$-wave pairings. The phase boundary between them is indicated by the red dashed lines.
@@
where $(1,2,3)\equiv(x,y,z)$. All the Hermitian matrices are represented as Kronecker products $\Gamma_{\mu\nu}=\tau_{\mu}\otimes\sigma_{\nu}$. Two sets of Pauli matrices $\tau_{\mu}$ and $\sigma_{\mu}$ act on the parity and spin indices, respectively, where $\mu=0,1,2,3$, and $\tau_{0}$ and $\sigma_{0}$ are the identity matrices. The associated spinor basis is $\Psi_{\boldsymbol{k}}^{\top}=[c_{+\uparrow},c_{+\downarrow},c_{-\uparrow},c_{-\downarrow}](\boldsymbol{k})$, where $c_{\tau\sigma}(\boldsymbol{k})$ is a fermion annihilation operator with parity $\tau=\pm$, spin $\sigma=\uparrow,\downarrow$, and momenta $\boldsymbol{k}$. Here, $v_{i}$ ($k_{i}$) is the Fermi velocity (momenta) along the $i$th coordinate and $\mu$ is the chemical potential. We consider a cubic system, for which $w_{x}=w_{y}=w_{z}\equiv w$ (say) for $w=v,a,b$. The Dirac Hamiltonian then enjoys the full O(3) rotational symmetry, generated by $\{\Gamma_{01},\Gamma_{02},\Gamma_{03}\}$. In Eq. (1), $m$ ($bk^{2}$) describes an isotropic constant (momentum-dependent Wilson-Dirac) mass, $\epsilon_{0}$ ($ak^{2}$) captures momentum independent (dependent) particle-hole asymmetry with $k^{2}\equiv|\boldsymbol{k}|^{2}$. For simplicity, we set $\epsilon_{0}=a=0$. Then the chemical potential is measured from the band center or zero energy. The massive Dirac Hamiltonian also possesses discrete parity ($\mathcal{P}$), time reversal ($\mathcal{T}$), and charge conjugation ($\mathcal{C}$) symmetries. In the announced basis, $\mathcal{P}\Psi_{\boldsymbol{k}}\mathcal{P}=\Gamma_{30}\Psi_{-\boldsymbol{k}}$, $\mathcal{T}\Psi^{\ast}_{\boldsymbol{k}}\mathcal{T}=-\Gamma_{02}\Psi_{-\boldsymbol{k}}$, and $\mathcal{C}\Psi_{\boldsymbol{k}}\mathcal{C}=-i\Gamma_{12}\Psi^{\ast}_{\boldsymbol{k}}$ [1, 38, 39, 40].
@@
The single-particle spectra of $\hat{h}_{\boldsymbol{k}}$, hosting two distinct phases for $\mu=0$, are ±$\pm\sqrt{v^{2}k^{2}+(m+bk^{2})^{2}}-\mu$. For $mb>0$, the system is in the topologically trivial phase with minimal spectral gap $2m$, and setting $\mu>m$ yields a simply connected FS. On the other hand, for $mb<0$, the system is in the topological phase. Inside the topological regime, for $12mb<-$ 1 the energy spectra show a “sagging” of the bands, thereby supporting an annular (simply connected) FS for $\sqrt{-1-4bm}/(2b)<\mu<m$ ($\mu>m$). By contrast, when $-1<2mb<0$, the FS in the doped TI is simply connected. From now on we set $b>0$ and tune the sign of $m$ to describe the trivial ($m>0$) and topological ($m<0$) phases. The band gap closes at a quantum critical point $m=0$, hosting massless Dirac excitations around $\boldsymbol{k}=0$.
@@
Interactions. In this work, we investigate the superconducting instabilities in doped Dirac insulators triggered by nonretarded local “Hubbardlike” interactions of the form $(\Psi^{\dagger}\Gamma_{\mu\nu}\Psi)(\Psi^{\dagger}\Gamma_{\rho\lambda}\Psi)$. Upon taking into account the cubic or rotational, $\mathcal{P}$, $\mathcal{T}$ and $\mathcal{C}$ symmetries, the interacting Lagrangian containing all (eight) such symmetry-allowed four-Fermion terms reads [39, 40]
@@
Here the superscript ${\rm s}$ (${\rm t}$) represents singlet (triplet) in the spin sector. But, the number of linearly independent quartic couplings is reduced to four via the Fierz identity [41], which we choose to be the ones in the spin-singlet channels. These derivations are shown in the Supplementary Material (SM) [42]. Then the imaginary time ($\tau$) action can be decomposed as $S=S_{0}+S_{\mathrm{int}}$, where
@@
and $\Psi^{\dagger}_{\tau,\boldsymbol{x}}$ and $\Psi_{\tau,\boldsymbol{x}}$ are independent Grassmann variables. Any microscopic model of local interactions can be captured by the vector coupling $\boldsymbol{g}^{s}_{\beta}$. Here we consider a multi-orbital Hubbard model with only on-site density-density repulsions ($U_{1},U_{2}>0$) with the Hamiltonian
@@
where $n_{\sigma}=n_{+\sigma}+n_{-\sigma}$ for $\sigma=\uparrow,\downarrow$, $n_{\tau}=n_{\tau\uparrow}+n_{\tau\downarrow}$ for $\tau=\pm$, and $n_{\tau\sigma}=c^{\dagger}_{\tau\sigma}c_{\tau\sigma}$ is the fermionic density with parity $\tau$ and spin projection $\sigma$. Single-orbital systems allow only $U_{1}$ (conventional on-site Hubbard repulsion), while $U_{2}$ is nontrivial due to the parity degrees of freedom. In terms of $\boldsymbol{g}^{s}_{\beta}$, the above model is captured by [42]
${}g_{0}^{s}=U_{1}/2+U_{2}/4,\>\>g_{1}^{s}=g_{2}^{s}=U_{1}/8,\>\>g_{3}^{s}=U_{1}/8-U_{2}/4.$ (6)
@@
Table 1: Local superconducting (excitonic) orders with their conjugate fields (CF, first column), matrices associated with the fermion bilinears $\Psi^{\dagger}_{\rm Nam}\eta_{\mu}\Gamma_{\nu\rho}\Psi_{\rm Nam}$ in the Nambu basis (second column), and their physical meaning (third column) are shown in the upper (lower) block. Here $j=1,2,3$ and in the superconducting channels $\alpha=1,2$, reflecting the U(1) gauge redundancy in the choice of the superconducting phase ($\phi$) [Eq. (Superconductivity in three-dimensional interacting doped topological insulators)]. Their transformations under the discrete ${\mathcal{P}}$, ${\mathcal{T}}$, and ${\mathcal{C}}$ symmetries are shown in the fourth, fifth and sixth columns, respectively, where $+$ (--) indicating even (odd).
Orders. For a systematic analysis of superconducting instabilities, we now introduce a Nambu spinor $\Psi_{\rm Nam}(\boldsymbol{k})=(\Psi_{\boldsymbol{k}},\Gamma_{02}\Psi^{\ast}_{-\boldsymbol{k}})$, where in the lower block we absorb the unitary part of the time reversal operator. The Dirac Hamiltonian in this basis reads as $\hat{h}_{\rm Nam}(\boldsymbol{k})=\eta_{3}\hat{h}_{\boldsymbol{k}}$, where $\eta_{\mu}$ is another set of Pauli matrices acting on the newly-introduced Nambu degree of freedom. The effective action containing all symmetry-allowed local pairing (pair) and excitonic (exc) orders takes the form
@@
The physical meanings and symmetry properties of all the local orders are summarized in Table 1. Around the FS, the (pseudo)scalar pairing takes the form of singlet $s$-wave (triplet $p$-wave) pairing around the FS [31].
@@
$\frac{\mathrm{d}\>{\rm ln}\Delta_{\beta}^{\alpha}}{\mathrm{d}\ell}-z=\sum_{\nu}F_{\nu}^{\beta}(t,\mu,m,b)g^{\rm s}_{\nu},$ (12)
with $F^{\beta}_{\nu}$ once again functions of $t$, $\mu$, $m$ and $b$. The functions $H_{\nu\rho}^{\beta}$ and $F_{\nu}^{\beta}$ are shown in the SM [42].
In the presence of an underlying FS, RG flows of $\boldsymbol{g}^{s}_{\beta}$, whose bare values at the scale of the ultraviolet cutoff $\Lambda$ or $\ell=0$ is set by Eq. (6), have to be stopped at an infrared scale $\ell^{\star}_{\mu}=-\ln(\mu(0))/z$, where $\mu(0)<1$ is the bare value of $\mu$ [43, 44, 45, 46]. In the disordered phase (without any symmetry breaking) all $g_{\beta}^{\rm s}$s decrease under coarse grain, and none of them diverge. Its phase boundary with the ordered phase is characterized by the divergence of at least one of the coupling constants at an RG scale $\ell_{\rm div}=\ell^{\star}_{\mu}$. Within the ordered phase the respective coupling constants diverge at an RG scale $\ell_{\rm div}<\ell^{\star}_{\mu}$. The exact nature of the leading instability or the pattern of symmetry breaking is then unambiguously identified by the renormalized susceptibility $\Delta_{\beta}^{\alpha}$ that receives the largest positive correction (anomalous dimension) when one of the coupling constants diverges at the RG scale $\ell_{\rm div}$.
Throughout, we set $t=10^{-6}$ to construct the zero temperature phase diagrams. Nonetheless, the notion of $\ell_{\rm div}$ provides a poor man’s estimation of the transition temperature ($t_{c}$) in the following way. The solution of the RG flow equation for dimensionless temperature from Eq. (11) reads as $t(\ell)=t(0)\exp(z\ell)$. In an interacting doped Dirac insulator, there are two infrared RG scales $\ell^{\star}_{\mu}$ and $\ell_{\rm div}$, respectively at which the renormalized chemical potential and at least one of the coupling constants becomes of the order of unity. Associated with these two RG scales, we can define two bare temperatures $t(0)=t_{\mu}$ and $t_{g}$, respectively, such that $t(\ell^{\star}_{\mu})=t(\ell_{\rm div})=1$, yielding $t_{\mu}=\exp[-z\ell^{\star}_{\mu}]$ and $t_{g}=\exp[-z\ell^{\star}_{\rm div}]$. The transition temperature in the ordered phase is then defined as $t_{c}=t_{g}-t_{\mu}$ (up to an overall numerical prefactor). Defined this way, $t_{c}=0$ at the phase boundary between the disordered and ordered (superconducting) phases, and inside the ordered phase $t_{c}$ increases monotonically with the increasing coupling strength, as then $\ell_{\rm div}$ decreases monotonically.
@@
$g_{0}^{s}$ $\Delta^{\rm p}_{0}$ O(2) SR2
$g_{1}^{s}$ $\Delta^{\rm s}_{1}$ O(1)/$Z_{2}$ SR1 $\Delta^{\rm p}_{0}$ O(2) SR2 O(3)
$g_{2}^{s}$ $\Delta^{\rm s}_{2}$ O(1)/$Z_{2}$ SR1 $\Delta^{\rm p}_{1}$ O(2) SR2 O(3)
$g_{3}^{s}$ $\Delta^{\rm s}_{3}$ O(1)/$Z_{2}$ SR1 $\Delta^{\rm p}_{0}$ O(2) SR2 O(3)
$g_{0}^{t}$ $\Delta^{\rm t}_{0}$ O(3) SR1 $\Delta^{\rm p}_{2}$ O(3) SR2 O(3)
$g_{1}^{t}$ $\Delta^{\rm t}_{1}$ O(3) SR1 $\Delta^{\rm p}_{3}$ O(2) SR2 O(5)
$g_{2}^{t}$ $\Delta^{\rm p}_{0}$ O(2) SR2
$g_{3}^{t}$ $\Delta^{\rm s}_{2}$ O(1)/$Z_{2}$ SR2 $\Delta^{\rm p}_{1}$ O(2) SR2 O(3)
@@
Phase diagram. We follow the outlined procedure to construct the phase diagram for the interacting doped 3D Dirac insulator in the $(U_{1},U_{2})$ plane for different underlying FS topologies [Fig. 1(a)]. The results are shown in Fig. 1(b)-(d). While repulsive $U_{1}$ favors nucleation of conventional scalar $s$-wave pairing, repulsive $U_{2}$ is conducive to the condensation for the odd-parity pseudoscalar $p$-wave pairing. As the relative strength of these two interactions is varied, there is a transition between them, around which, in principle, the system can also foster an axionic $p+is$ paired state [30], which, however, cannot be captured from the RG analysis. These outcomes are insensitive to the underlying FS topology.

### v2

@@ Abstract
Three-dimensional doped Dirac insulators foster simply connected (in both topological and trivial regimes) and annular (deep inside the topological regime) Fermi surfaces (FSs) in the normal state, and allow on-site repulsions among fermions with opposite spin ($U_{1}$) and parity ($U_{2}$) eigenvalues. From an unbiased leading-order (one-loop) renormalization group analysis, controlled by a suitable $\epsilon$ expansion, we show that this system develops a strong propensity toward the nucleation of scalar $s$-wave and odd-parity pseudoscalar $p$-wave pairings, favored by repulsive $U_{1}$ and $U_{2}$ interactions, respectively, irrespective of the underlying FS topology. Our results can be pertinent for the observed superconductivity in various doped narrow gap semiconductors, and the theoretical foundation can readily be applied to investigate similar phenomenon in various doped topological materials.
@@
Introduction. From the baryonic universe to topological quantum crystals, Dirac theory offers a unified description. As such electrical and thermal topological insulators (TIs) belonging to any Altland-Zirnbauer symmetry class in any dimension ($d$) can be modeled by a suitable massive Dirac Hamiltonian [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The ${\bf k}\cdot{\bf p}$ Hamiltonian for their candidate materials also assumes a similar form in the low energy and long-wavelength limit [16]. Fascinatingly, three-dimensional (3D) doped or intercalated TIs, (Cu/Nd/Sr)xBi2Se3 and Sn1-xInxTe, with a Fermi surface (FS), become a fully gapped superconductor at the low temperatures [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The paired state is possibly topological ($p$-wave) in nature [27, 28, 29, 30, 31], showing a zero bias conductance peak (ZBCP), stemming from surface Majorana fermions. It is also conceivable to induce a transition between topological $p$-wave and trivial $s$-wave (devoid of ZBCP) pairings in Sn1-xInxTe. Despite such exciting experimental findings in the last decade, the phase diagram of interacting doped TIs, fostering superconducting ground states still remains hazy.
@@
Doped Dirac insulators accommodate simply connected and annular FSs, see Fig. 1(a). In this Letter, we theoretically study the effects of electron-electron ($e$-$e$) interactions on the phase diagram of such systems using an unbiased leading-order renormalization group (RG) analysis, controlled by an appropriate $\epsilon$ expansion. In strongly correlated TIs, such as the ones in Kondo systems [32, 33, 34], $e$-$e$ interactions can arise from Hubbardlike repulsions. In weakly correlated materials, effective $e$-$e$ interactions can be mediated by optical phonons below the scale of optical frequency [35, 36, 37]. Without delving into their exact microscopic origin, here we consider all symmetry allowed $e$-$e$ interactions. Then within the framework of a Hubbardlike model, composed of the on-site repulsions among fermions with opposite spin ($U_{1}$) and parity ($U_{2}$) eigenvalues, we show that 3D correlated doped TIs feature an intriguing confluence of odd-parity pseudoscalar topological pairing and trivial scalar $s$-wave pairing, respectively favored by $U_{2}$ and $U_{1}$, irrespective of the underlying FS geometry. See Figs. 1(b)-(d).
@@
Figure 1: (a) Fermi surface (FS) geometries for various chemical doping (dashed colored lines) in a Dirac insulator in the topological ($m<0$) and trivial ($m>0$) regimes with $b>0$. A quantum phase transition between them takes place at $m=0$, around which the FS is simply connected (red spheres). Deep inside the topological regime ($m<-1/(2b)$), the system supports annular FS (concentric blue spheres). Black curves show energy eigenvalues (vertical direction) along a momentum (horizontal direction). Phase diagram in the $(U_{1},U_{2})$ plane for (b) $\mu=0.685$, $m=0.7$, $b=1.0$ (yielding an annular FS in doped TI), (c) $\mu=0.730$, $m=-0.7$, $b=1.0$ (yielding a simply connected FS in doped TI), (d) $\mu=0.180$, $m=0.1$, $b=0.1$ (yielding a simply connected FS in doped trivial insulator) and $t=10^{-6}$ (mimics zero temperature). Color map in (b)-(d) indicates the transition temperature ($t_{c}$). White regions mark a Fermi liquid without any ordering. The phase diagrams display odd-parity topological $p$-wave and trivial $s$-wave pairings, with the red dashed lines indicating the phase boundary between them.
@@
where $(1,2,3)\equiv(x,y,z)$. All the Hermitian matrices are represented as Kronecker products $\Gamma_{\mu\nu}=\tau_{\mu}\otimes\sigma_{\nu}$. Two sets of Pauli matrices $\tau_{\mu}$ and $\sigma_{\mu}$ act on the parity and spin indices, respectively, where $\mu=0,1,2,3$, and $\tau_{0}$ and $\sigma_{0}$ are the identity matrices. The associated spinor basis is $\Psi_{\boldsymbol{k}}^{\top}=[c_{+\uparrow},c_{+\downarrow},c_{-\uparrow},c_{-\downarrow}](\boldsymbol{k})$, where $c_{\tau\sigma}(\boldsymbol{k})$ is a fermion annihilation operator with parity $\tau=\pm$, spin $\sigma=\uparrow,\downarrow$, and momenta $\boldsymbol{k}$. Here, $v_{i}$ ($k_{i}$) is the Fermi velocity (momenta) along the $i$th coordinate and $\mu$ is the chemical potential. We consider a cubic system, for which $w_{x}=w_{y}=w_{z}\equiv w$ (say) for $w=v,a,b$. The Dirac Hamiltonian then enjoys the full $O(3)$ rotational symmetry, generated by $\{\Gamma_{01},\Gamma_{02},\Gamma_{03}\}$. In Eq. (1), $m$ ($bk^{2}$) describes an isotropic constant (momentum-dependent Wilson-Dirac) mass, $\epsilon_{0}$ ($ak^{2}$) captures momentum-independent (momentum-dependent) particle-hole asymmetry with $k^{2}\equiv|\boldsymbol{k}|^{2}$. For simplicity, we set $\epsilon_{0}=a=0$. Then the chemical potential is measured from the band center or zero energy. The massive Dirac Hamiltonian also possesses discrete parity ($\mathcal{P}$), time reversal ($\mathcal{T}$), and charge conjugation ($\mathcal{C}$) symmetries. In the announced basis, $\mathcal{P}\Psi_{\boldsymbol{k}}\mathcal{P}=\Gamma_{30}\Psi_{-\boldsymbol{k}}$, $\mathcal{T}\Psi^{\ast}_{\boldsymbol{k}}\mathcal{T}=-\Gamma_{02}\Psi_{-\boldsymbol{k}}$, and $\mathcal{C}\Psi_{\boldsymbol{k}}\mathcal{C}=-i\Gamma_{12}\Psi^{\ast}_{\boldsymbol{k}}$ [1, 38, 39, 40].
@@
The single-particle spectra of $\hat{h}_{\boldsymbol{k}}$, hosting two distinct phases for $\mu=0$, are ±$\pm\sqrt{v^{2}k^{2}+(m+bk^{2})^{2}}-\mu$. For $mb>0$, the system is in the trivial phase with minimal spectral gap $2m$, and setting $\mu>m$ yields a simply connected FS. On the other hand, for $mb<0$, the system is in the topological phase. Inside the topological regime, for $12mb<-$ 1 the energy spectra show a “sagging” of the bands, thereby supporting an annular (simply connected) FS for $\sqrt{-1-4bm}/(2b)<\mu<m$ ($\mu>m$). By contrast, when $-1<2mb<0$, the FS in the doped TI is simply connected. From now on we set $b>0$ and tune the sign of $m$ to describe the trivial ($m>0$) and topological ($m<0$) phases. The band gap closes at a quantum critical point $m=0$, hosting massless Dirac excitations around $\boldsymbol{k}=0$.
@@
Interactions. In this Letter, we investigate the superconducting instabilities in doped Dirac insulators triggered by nonretarded local “Hubbardlike” interactions of the form $(\Psi^{\dagger}\Gamma_{\mu\nu}\Psi)(\Psi^{\dagger}\Gamma_{\rho\lambda}\Psi)$. Upon taking into account the cubic or rotational, $\mathcal{P}$, $\mathcal{T}$ and $\mathcal{C}$ symmetries, the interacting Lagrangian containing all (eight) such symmetry-allowed four-fermion terms reads [39, 40]
@@
Here, the superscript ${\rm s}$ (${\rm t}$) represents a singlet (triplet) in the spin sector. But, the number of linearly independent quartic couplings is reduced to four via the Fierz identity [41], which we choose to be the ones in the spin-singlet channels. These derivations are shown in the Supplemental Material (SM) [42]. Then the imaginary time ($\tau$) action can be decomposed as $S=S_{0}+S_{\mathrm{int}}$, where
@@
and $\Psi^{\dagger}_{\tau,\boldsymbol{x}}$ and $\Psi_{\tau,\boldsymbol{x}}$ are independent Grassmann variables. Any microscopic model of local interactions can be captured by the vector coupling $\boldsymbol{g}^{\rm s}$. Here we consider a multi-orbital Hubbard model with only on-site density-density repulsions ($U_{1},U_{2}>0$) with the Hamiltonian
@@
where $n_{\sigma}=n_{+\sigma}+n_{-\sigma}$ for $\sigma=\uparrow,\downarrow$, $n_{\tau}=n_{\tau\uparrow}+n_{\tau\downarrow}$ for $\tau=\pm$, and $n_{\tau\sigma}=c^{\dagger}_{\tau\sigma}c_{\tau\sigma}$ is the fermionic density with parity $\tau$ and spin projection $\sigma$. Single-orbital systems allow only $U_{1}$ (conventional on-site Hubbard repulsion), while $U_{2}$ is nontrivial due to the parity degrees of freedom. In terms of $\boldsymbol{g}^{\rm s}$, the above model is captured by [42]
${}g_{0}^{\rm s}=U_{1}/2+U_{2}/4,\>\>g_{1}^{\rm s}=g_{2}^{s}=U_{1}/8,\>\>g_{3}^{\rm s}=U_{1}/8-U_{2}/4.$ (6)
@@
Table 1: Local superconducting (excitonic) orders with their conjugate fields (CF, first column), matrices associated with the fermion bilinears $\Psi^{\dagger}_{\rm Nam}\eta_{\mu}\Gamma_{\nu\rho}\Psi_{\rm Nam}$ in the Nambu basis (second column), and their physical meaning (third column) are shown in the upper (lower) block. Here, $j=1,2,3$ and in the superconducting channels $\alpha=1,2$, reflecting the $U(1)$ gauge redundancy in the choice of the superconducting phase ($\phi$) [Eq. (Superconductivity in three-dimensional interacting doped topological insulators)]. Their transformations under the discrete ${\mathcal{P}}$, ${\mathcal{T}}$, and ${\mathcal{C}}$ symmetries are shown in the fourth, fifth and sixth columns, respectively, where $+$ (--) indicating even (odd).
Orders. For a systematic analysis of superconducting instabilities, we now introduce a Nambu spinor $\Psi_{\rm Nam}(\boldsymbol{k})=(\Psi_{\boldsymbol{k}},\Gamma_{02}\Psi^{\ast}_{-\boldsymbol{k}})$, where in the lower block we absorb the unitary part of the time-reversal operator. The Dirac Hamiltonian in this basis reads as $\hat{h}_{\rm Nam}(\boldsymbol{k})=\eta_{3}\hat{h}_{\boldsymbol{k}}$, where $\eta_{\mu}$ is another set of Pauli matrices acting on the newly introduced Nambu degree of freedom. The effective action containing all symmetry-allowed local pairing (pair) and excitonic (exc) orders takes the form
@@
The physical meanings and symmetry properties of all the local orders are summarized in Table 1. Around the FS, the scalar (pseudoscalar) pairing takes the form of singlet $s$-wave (triplet $p$-wave) pairing [31].
@@
$\frac{\mathrm{d}\>{\rm ln}\Delta_{\beta}^{\alpha}}{\mathrm{d}\ell}-z=\sum_{\nu}F_{\nu}^{\beta}(t,\mu,m,b)g^{\rm s}_{\nu}.$ (12)
The functions $H_{\nu\rho}^{\beta}$ and $F_{\nu}^{\beta}$ are shown in the SM [42].
In the presence of an underlying FS, RG flows of $\boldsymbol{g}^{\rm s}_{\beta}$, whose bare values at the scale of the ultraviolet cutoff $\Lambda$ or $\ell=0$ is set by Eq. (6), have to be stopped at an infrared scale $\ell^{\star}_{\mu}=-\ln(\mu(0))/z$, where $\mu(0)<1$ is the bare value of $\mu$ [43, 44, 45, 46]. In the disordered phase (without any symmetry breaking) all $g_{\beta}^{\rm s}$s decrease under coarse grain, and none of them diverge. Its phase boundary with the ordered phase is characterized by the divergence of at least one of the coupling constants at an RG scale $\ell_{\rm div}=\ell^{\star}_{\mu}$. Within the ordered phase the respective coupling constants diverge at an RG scale $\ell_{\rm div}<\ell^{\star}_{\mu}$. The exact nature of the leading instability or the pattern of symmetry breaking is then unambiguously identified by the renormalized susceptibility $\Delta_{\beta}^{\alpha}$ that receives the largest positive correction (anomalous dimension) when one of the coupling constants diverges at the RG scale $\ell_{\rm div}$.
Throughout, we set $t=10^{-6}$ to construct the zero-temperature phase diagrams. Nonetheless, the notion of $\ell_{\rm div}$ provides a poor man’s estimation of the transition temperature ($t_{c}$) in the following way. The solution of the RG flow equation for dimensionless temperature from Eq. (11) reads as $t(\ell)=t(0)\exp(z\ell)$. In an interacting doped Dirac insulator, there are two infrared RG scales $\ell^{\star}_{\mu}$ and $\ell_{\rm div}$, respectively at which the renormalized chemical potential and at least one of the coupling constants becomes of the order of unity. Associated with these two RG scales, we can define two bare temperatures $t(0)=t_{\mu}$ and $t_{g}$, respectively, such that $t(\ell^{\star}_{\mu})=t(\ell_{\rm div})=1$, yielding $t_{\mu}=\exp[-z\ell^{\star}_{\mu}]$ and $t_{g}=\exp[-z\ell^{\star}_{\rm div}]$. The transition temperature in the ordered phase is then defined as $t_{c}=t_{g}-t_{\mu}$ (up to an overall numerical prefactor). Defined this way, $t_{c}=0$ at the phase boundary between the disordered and ordered (superconducting) phases, and inside the ordered phase $t_{c}$ increases monotonically with the increasing coupling strength, as then $\ell_{\rm div}$ decreases monotonically.
@@
$g_{0}^{\rm s}$ $\Delta^{\rm p}_{0}$ $O(2)$ SR2
$g_{1}^{\rm s}$ $\Delta^{\rm s}_{1}$ $O(1)$/$Z_{2}$ SR1 $\Delta^{\rm p}_{0}$ $O(2)$ SR2 $O(3)$
$g_{2}^{\rm s}$ $\Delta^{\rm s}_{2}$ $O(1)$/$Z_{2}$ SR1 $\Delta^{\rm p}_{1}$ $O(2)$ SR2 $O(3)$
$g_{3}^{\rm s}$ $\Delta^{\rm s}_{3}$ $O(1)$/$Z_{2}$ SR1 $\Delta^{\rm p}_{0}$ $O(2)$ SR2 $O(3)$
$g_{0}^{\rm t}$ $\Delta^{\rm t}_{0}$ $O(3)$ SR1 $\Delta^{\rm p}_{2}$ $O(3)$ SR2 $O(3)$
$g_{1}^{\rm t}$ $\Delta^{\rm t}_{1}$ $O(3)$ SR1 $\Delta^{\rm p}_{3}$ $O(2)$ SR2 $O(5)$
$g_{2}^{\rm t}$ $\Delta^{\rm p}_{0}$ $O(2)$ SR2
$g_{3}^{\rm t}$ $\Delta^{\rm s}_{2}$ $O(1)$/$Z_{2}$ SR2 $\Delta^{\rm p}_{1}$ $O(2)$ SR2 $O(3)$
@@
Phase diagram. We follow the outlined procedure to construct the phase diagram for the interacting doped 3D Dirac insulator in the $(U_{1},U_{2})$ plane for different underlying FS topologies [Fig. 1(a)]. The results are shown in Figs. 1(b)-(d). While repulsive $U_{1}$ favors nucleation of conventional scalar $s$-wave pairing, repulsive $U_{2}$ is conducive to the condensation for the odd-parity pseudoscalar $p$-wave pairing. As the relative strength of these two interactions is varied, there is a transition between them, around which, in principle, the system can also foster an axionic $p+is$ paired state [30], which, however, cannot be captured from the RG analysis. These outcomes are insensitive to the underlying FS topology.